8532964573:

Solve the equation using the distributive property and properties of equality. One-half (x + 6) = 18 What is the value of x? 6 7 and one-half 14 and one-half 30

jhonyy9:

\[\frac{ 1 }{ 2 }(x+6) = 18 \] so like a first step multiplie both sides by 2 ,what you get what will can you simplifie in this way ?

supie:

Simplify: \(\LARGE\frac{1}{2}(x + 6) = 18\) Distribute: \(\LARGE(\frac12)(x)+(\frac12)(6)\) \(\LARGE =8\) -3, both sides: \(\LARGE\frac12x+3−3=18−3\) \(\LARGE\frac12x=15\) *2, both sides: \(\LARGE2*(\frac{1}{2}x)=(2)*(15)\)